# Evidence for large baryonic feedback at low and intermediate redshifts from kinematic Sunyaev-Zel’dovich observations with ACT and DESI photometric galaxies

### v1

@@ Abstract
Recent advances in cosmological observations have provided an unprecedented opportunity to investigate the distribution of baryons relative to the underlying matter. In this work, we robustly show that the gas is much more extended than the dark matter at 40 and the amount of baryonic feedback at strongly disfavors low-feedback models such as that of state-of-the-art hydrodynamical simulation IllustrisTNG compared with high-feedback models such as that of the original Illustris simulation. This has important implications for bridging the gap between theory and observations and understanding galaxy formation and evolution. Furthermore, a better grasp of the baryon-dark matter link is critical to future cosmological analyses, which are currently impeded by our limited knowledge of baryonic feedback. Here, we measure the kinematic Sunyaev-Zel’dovich (kSZ) effect from the Atacama Cosmology Telescope (ACT), stacked on the luminous red galaxy (LRG) sample of the Dark Energy Spectroscopic Instrument (DESI) imaging survey. This is the first analysis to use photometric redshifts for reconstructing galaxy velocities. Due to the large number of galaxies comprising the DESI imaging survey, this is the highest signal-to-noise stacked kSZ measurement to date: we detect the signal at 13 and find that the gas is more spread out than the dark matter at 40. Our work opens up the possibility to recalibrate large hydrodynamical simulations using the kSZ effect. In addition, our findings point towards a way of alleviating inconsistencies between weak lensing surveys and cosmic microwave background (CMB) experiments such as the ‘low ’ tension, and shed light on long-standing enigmas in astrophysics such as the ‘missing baryon’ problem.
@@ I Introduction
In this work, we measure the baryon profiles around Dark Energy Spectroscopic Instrument (DESI) luminous red galaxies (LRGs) using the stacking of the kSZ effect, which has been measured using a variety of techniques over the years [9, 10, 11, 12, 13]. Our findings shed light on the complex relationship between baryonic matter and dark matter. The novelty of this work lies in the following three aspects: 1) its pioneering use of photometric redshifts for probing the stacked kSZ effect, which departs from the traditional use of spectroscopic redshifts; 2) challenging the predictions of state-of-the-art hydrodynamical simulations – we find that the feedback processes appear to be stronger in the real Universe than in simulations – in other words, the baryons are expelled more violently from the cores of galaxy groups and clusters; 3) its implications for alleviating or resolving discrepancies in weak lensing analysis; most notably, the ‘low ’ tension found when comparing the inferred values of the clumpiness parameter, , from weak lensing with CMB probes (e.g., see [14] and references therein), and the ‘Lensing is low’ tension, which constitutes a discrepancy between galaxy clustering and galaxy-galaxy lensing [15, 16].
@@ II.1 DESI imaging survey
The Dark Energy Spectroscopic Instrument is a robotic, fiber-fed, highly multiplexed spectroscopic surveyor that operates on the Mayall 4-meter telescope at Kitt Peak National Observatory [17]. DESI, which can obtain simultaneous spectra of almost 5000 objects over a field [18, 19, 20], is currently conducting a five-year dark energy survey of about a third of the sky [21]. This campaign will obtain spectra for approximately 40 million galaxies and quasars [22].
In this paper, we use the photometric sample of DESI Luminous Red Galaxies (LRGs) and a higher-density extended LRG sample, referred to as ‘Main LRGs’ and ‘Extended LRGs’ [23, 24, 25, 26, 27]. The parent imaging is the DESI Legacy Imaging Survey, which is a combination of optical and mid-infrared imaging used for DESI target selection from three telescopes: Blanco for Dark Energy Camera Legacy Survey (DECaLS), Mayall for the Mayall -band Legacy Survey (MzLS), and Bok for the Beijing–Arizona Sky Survey (BASS). We also make use of the reliable photometric redshifts presented in Zhou et al. [27] for Data Release 9 (DR9) and Data Release 10 (DR10), which have been calibrated via DESI spectroscopic redshifts and thus have only a small fraction of contaminants and a mean error of . Similarly to Zhou et al. [27], we define four tomographic bins: [0.4, 0.54], [0.54 0.713], [0.713, 0.86] and [0.86 1.024] (see Table 1 for mean bin redshifts and number of galaxies in the overlap with ACT). We take into account the estimated error in the photometric redshifts for each galaxy, by imposing a maximum cut on the estimated error of . This removes 3% of the galaxies, including outliers with anomalously large redshift errors that could otherwise bias the reconstructed velocities by diluting the density signal [see 28, for tests on simulations]. In App. B, we investigate the effects of not imposing such a cut and find that, for the most part, they are negligible.
@@ II.2 ACT temperature maps
This paper employs the harmonic-space Internal Linear Combination (hILC) maps [29, 30] from the Data Release 6 (DR6) from the Atacama Cosmology Telescope (ACT), a 6 m telescope that was located in the Atacama Desert of Chile and measured the CMB from 2007 to 2022. The DR6 data comprise multifrequency observations from 2017 to 2022 and the hILC map covers roughly a third of the sky (32%) at three frequency bands: f090 (77–112 GHz), f150 (124–172 GHz), and f220 (182–277 GHz), and uses Planck data on large scales [31, 32]. The observational program of DR6 targeted the ‘wide’ field. For this work, we use only the night-time portion of the data taken in the first five observing seasons (2017-2021). The ACT maps are produced in the plate-carrée (CAR) projection scheme. This analysis uses the first version of the ACT DR6 maps, dr6.01. Similarly to [33], we apply a mask on the ACT map that removes all detected clusters [34] and point sources as well as outliers in any of the filtered temperature bins, since those can bias the stacked profiles.
@@ III.1 Reconstruction
In order to obtain an estimate of the peculiar velocity of each galaxy, we adopt the standard BAO reconstruction method of [36], implemented in the package pyrecon 111https://github.com/cosmodesi/pyrecon [37]. This method yields an estimate of the first-order galaxy displacement field, which can be converted into an estimate for the velocity. A study performed on realistic DESI-like light cone mocks [28] informs us that the cross-correlation coefficient, , between the reconstructed galaxy velocities along the line-of-sight (denoted by subscript ) and the host halo velocities, which captures imperfections in the velocity reconstruction, is about (with an uncertainty of about 10%), the value we adopt in this study, while that of the spectroscopic sample is about . In other words, the reconstruction performance deteriorates by about half when using photometric redshifts, a prediction also confirmed by a similar study in the ‘snapshot geometry’ [38]. Nonetheless, the measurement around the photometric LRG sample yields a high signal-to-noise ratio (SNR) due to the large number of galaxies comprising the imaging survey. We note that LRGs are found mostly in galaxy groups with a mean linear bias of [39]. In addition to the cuts in [33], for our fiducial analysis, we remove outliers in the reconstructed velocities and CAP-filtered temperature at the largest apertures measured (6 arcmin) at 3222Objects that are simultaneous outliers in both temperature and reconstructed velocity at 3 constitute 0.1%. and ensure that the number of galaxies in each velocity bin is symmetric around the mean by random downsampling, which avoids unwanted biases from massive clusters and guarantees that in the absence of a kSZ signal, our estimator yields zero mean signal. In App. B, we show that these corrections have a mostly negligible effect: the uncorrected measurements tend to have marginally higher SNR (as a result of the larger number of galaxies in the samples) except for bin 3, which has a lower SNR due to its slightly asymmetric line-of-sight velocity distribution that improves once we remove the outliers.
@@ III.2 Estimator
Once we have estimates of the velocities, we can measure the CMB temperatures around each galaxy using a compensated aperture photometry (CAP) filter [33]. There are several benefits to using the CAP filter. Unlike a matched filter, the CAP filter is agnostic about the profile shape and because it is measured at different radii, it allows us to reconstruct the spherically averaged profiles. It is very effective at removing the large-scale primary CMB, as well as the uncorrelated part of the signal, leaving, in principle, only the sum of the one- and two-halo terms [see 40, for tests on simulations]. Additionally, the CAP-filtered profiles behave similarly to a cumulative density profile for large radii. The CAP filter is defined as:
@@
where is the rms of the radial component of the reconstructed velocities, , is the speed of light, and is the cross-correlation coefficient between the reconstructed and true velocity, evaluated from mock simulations. We use the publicly available pipeline ThumbStack333https://github.com/EmmanuelSchaan/ThumbStack to apply the estimator to the DESI and ACT data.
To interpret the kSZ profiles obtained through the above method, we need to estimate the covariance of the stacked measurements. We do so by applying a bootstrap resampling to the signal at each galaxy location. In particular, we draw 10,000 realizations of the galaxy catalogs (with repetition), and infer the covariance matrices from the scatter across the resampled stacked profiles (see Section IV.5).
@@ IV.1 Stacked kSZ signal per redshift bin
| LRG sample | # of galaxies | ||||
| Ext. DR9 | 963,631 | 0.47 | 43.2 | 19.9 | 6.4 |
| Ext. DR9 | 1,658,313 | 0.63 | 69.4 | 21.6 | 7.1 |
| Ext. DR10 | 1,951,646 | 0.79 | 80.3 | 26.1 | 8.2 |
| Ext. DR10 | 1,690,171 | 0.92 | 34.0 | 14.6 | 4.6 |
| Ext. DR9+10 all | 6,850,072 | 0.75 | 203.0 | 42.9 | 13.5 |
| Main DR9 | 422,350 | 0.47 | 25.2 | 13.7 | 4.4 |
| Main DR9 | 795,393 | 0.63 | 78.8 | 24.4 | 7.8 |
| Main DR10 | 753,945 | 0.79 | 65.9 | 23.4 | 7.3 |
| Main DR10 | 629,367 | 0.93 | 20.8 | 10.1 | 3.2 |
| Main DR9+10 all | 2,882,904 | 0.74 | 166.7 | 38.8 | 12.1 |
Fig. 1 shows the stacked kSZ signal (Eq. 4) from ACT for the four photometric bins of the DR10 Extended and Main LRG samples obtained from the DESI Imaging Survey (see Section II.1). The signal is shown as a function of CAP filter radius and detected at a high significance. We note that the error bars at large aperture are highly correlated (see Fig. 5). To aid interpretation, we convert the and axis into proper distance and CAP-filtered optical depth (), respectively, calculated at the mean redshift, . We note that the optical depth measures the integrated gas density along the line-of-sight,
@@
In Table 1, we show the significance of detection and corresponding SNR and chi-squared values for each of the four bins as well as the combination of all four bins. We define the SNR with respect to null444Corresponding to the inverse of the fractional error on the fit of a single amplitude parameter to the data. This often called the “detection significance” in number of Gaussian standard deviations . as , with best-fit coming from the Illustris-1 simulation555https://www.tng-project.org/data/ curve at (see Fig. 2 for more details) with a freed up parameter for the amplitude, and the SNR with respect to the dark matter distribution (with amplitude fixed by the theory best fit), defined as . We detect the signal at 13 and find the profiles to differ from the dark matter ones by . The dark matter profiles are obtained by stacking on dark-matter-only maps computed using the TNG300-1-Dark simulation (dark-matter-only counterpart to TNG300-1), and we have tested that other dark-matter-only simulations (e.g., Illustris-1-Dark) yield fully consistent results, as expected for pure -body simulations. The metric is defined in the standard way using the covariance matrix from Fig. 5:
where is our stacked kSZ measurementfrom the data, and is the model we are comparing against (if null, ).
To take advantage of the larger number of objects in DR9 and their relatively smaller photometric errors at low (compared with high ), we quote the DR9 results for bin 1 and 2 and the DR10 results for bin 3 and 4. For the combined ‘DR9+10 all,’ we use DR10 reconstructed velocities where available and otherwise DR9 ones in order to maximize the signal. We provide detailed comparison between DR9 and DR10 in Tables 2 and 3 finding a high level of congruence between the two. Compared with the previous measurement using CMASS LRGs and ACT DR4 [33], (same number of radial bins), we see that the total chi-squared for our Extended and Main samples is and 170, respectively, or about twice higher. In App. A, we find our curves to be in excellent agreement with [33].
Intuitively, one can think of these curves as showing roughly the cumulative gas density distribution. At large radii, we expect the profiles to become shallower as more and more of the gas is enclosed. We see that the gas profiles are steep beyond the virial radius, suggesting that a significant fraction of the gas resides beyond it (see discussion of Fig. 2, where we test this conjecture). The decline in amplitude at large apertures of bin 3 is due to noise from the primary CMB and the larger fraction of photometric outliers in that bin (see App. B).
@@ IV.2 Comparison with simulations
In Fig. 2, we study the comparison between the state-of-the-art hydrodynamical simulations IllustrisTNG-300-1 (TNG300-1) [41] and the measured gas profiles from the data. We also show a comparison with the older model of the Illustris-1 simulation [42], which has known shortcomings with predicting observations such as galaxy morphologies and colors, various cluster properties, and gas fractions [43, 44, 41]. We select LRG-like galaxies in the two simulations via an abundance matching approach and stack them at their locations on 2D CMB data mimicking ACT maps666Prior to applying any filter, these maps are convolved with the ACT beam, in this case a Gaussian with FWHM = 1.6 arcmin before being passed to the stacking pipeline.. Namely, we rank-order the galaxies by stellar mass and select the top such that, for the Main sample and for the Extended sample. We find that the mean halo mass of LRGs in TNG300-1 matches very well the inferred mean halo mass of DESI LRGs at , , and at , [39]. In Section IV.4, we split the galaxies into stellar mass bins estimated in [26] and detect the mass evolution of the signal at high significance.
Of most interest in this study is comparing the shapes of the profiles rather than the overall normalization, as the shapes encode information about the extent of the gas and hence the strength of the baryonic feedback effects. To make the comparison more straightforward, we multiply the TNG300-1 and Illustris-1 curves by a factor to match the amplitude of the data at the penultimate CAP radial bin. We find that there is a noticeable gap between the TNG300-1 prediction and the observed profiles, whereas the Illustris-1 model appears to be in much better agreement with the data. This result suggests stronger feedback is needed than present in TNG300-1, of similar magnitude to the older Illustris-1 model, which is known to have strong baryonic feedback. This result also suggests that the ‘ tension’ in weak lensing [14] could be partially explained by baryonic effects (see also Ref. [45]), and we leave a more detailed study for future work.
@@
We perform a number of tests on both hydrodynamical simulations to ensure that the gap we see is significant: we vary the satellite fraction from 0% to 30%777Centrals are selected as the most massive subhalo in each host halo. To vary the satellite fraction, we randomly discard centrals or satellites. (note that the constraints from [39] are ), we put all galaxies at the centers of their host halos, we vary the number density (between half and twice the fiducial value, i.e. from to ) and thus stellar mass threshold of the extracted LRGs, and we add noise to the halo velocities. In Fig. 2, we show the default scenario for both simulations and in shaded color for TNG300-1, we display the minimum and maximum deviations from the default caused by all the aforementioned alternative scenarios. Despite considering scenarios that push the bounds of physically reasonable models, we see that the gap between simulations and observations persists. In fact, we find a significant deviation of 17 (19) between the kSZ signal of the combined Main (Extended) sample and that of TNG300-1, defined as where once again we adopt the Illustris curve at as our theory curve with a free amplitude that is then used to rescale the TNG300-1 curve by. We stress that in quoting these numbers we fully marginalize over the amplitude of the signal. Future simulation efforts would need to focus on reconciling their outputs with our observed gas profiles from kSZ.
IV.3 Null test
In Fig. 3, we validate that our measured signal is indeed due to the kSZ effect rather than correlated contaminants such as the cosmic infrared background (CIB) or tSZ. For this test, we shuffle randomly the sample of reconstructed velocities to obtain 1000 realizations and perform the stacking of Eq. 4 but using the shuffled velocities on each galaxy. Since the true kSZ effect is proportional to the line-of-sight velocity, deleting that information in the shuffling should lead to a null measurement, which is indeed what we find. The tSZ and CIB effects do not have such dependence888Foreground emission such as from the CIB is also Doppler-boosted and acquires a term proportional to velocity. However, we find that this contribution is about two orders of magnitude smaller than the kSZ signal [46], and we’ll therefore neglect it in this analysis., and hence if our ACT map is correctly cleaned, any additive contamination should also vanish. We note that this argument also holds true for the unshuffled (original) kSZ measurement, which is robust to additive contributions.
IV.4 Mass evolution
Understanding the mass evolution of baryonic feedback is key to properly modeling its effect on cosmological observables such as cosmic shear, and the SZ effect is a versatile probe capable of capturing multiple gas properties [see e.g., 47, 48, 49, 50, 51, 52, for novel probes and methods]. Here we use estimates of the stellar mass from Zhou et al. [26] to split the galaxies in the combined (DR9+DR10) Main sample into four mass/luminosity bins: , each containing 926,232, 1,362,717, 458,886, and 2,152 galaxies, respectively. We note that the estimated masses have not been thoroughly characterized, and thus, they are mostly used to make qualitative statements.
We show the kSZ stacked measurements for each bin in Fig. 4. We detect the individual signals at high significance. The null is , 68.2, 96.2, and 28.3, and the deviation from the dark-matter profiles is , 24, 27, and 15, for each bin, respectively. As can be seen in the figure, we find a strong signature of the evolution of the amplitude with mass: namely, it increases with stellar mass/luminosity, as the optical depth is proportional to the total mass of the halo, i.e., the amplitudes of the stellar-mass bin samples differ roughly by the ratios of their respective mean stellar masses, since . We note that the relative differences between the amplitudes are independent of the calibration of the cross-correlation coefficient and the rms of the velocity. In addition, as expected, the curves become shallower at small apertures, as the mass increases. This is the case because massive halos retain more gas within the virial radius, as they have deeper potential wells compared with their less massive counterparts. At large apertures, the points are extremely strongly correlated, and it is imprudent to draw any definite conclusions on this issue (see Fig. IV.5). We leave a detailed analysis of the effect of baryons on weak lensing probes, as measured in this paper, for future work.
IV.5 Covariance matrix
The covariance matrix of the data is obtained by drawing 10,000 realizations of the galaxy catalogs (with repetition) and computing the covariance matrices from the resampled stacked profiles, i.e. bootstrapping. This produces an unbiased estimate of the covariance, in the limit of independent noise realizations from galaxy to galaxy. The assumption is not correct for large apertures where the temperature map cutouts (submaps) overlap, and the inferred covariance is only accurate at the 10% level (see [33]). However, given our errorbars, this level of accuracy is sufficient.
In Fig. 6, we illustrate the stacked 2D kSZ signal for the Main DR9 sample of DESI LRGs. We apply a Wiener filter () that effectively high-pass filters the CMB temperature map to isolate the small-scale signal that is the most affected by the kSZ effect. We can see by eye the extended gas envelope of the DESI LRGs, which spans several arcmin. This roughly corresponds to the mean halo virial radius at that redshift.

### v2

@@ Abstract
Recent advances in cosmological observations have provided an unprecedented opportunity to investigate the distribution of baryons relative to the underlying matter. In this work, we show that the gas is more extended than the dark matter, and the amount of baryonic feedback at disfavors low-feedback models such as that of state-of-the-art hydrodynamical simulation IllustrisTNG compared with high-feedback models such as that of the original Illustris simulation. This has important implications for bridging the gap between theory and observations and understanding galaxy formation and evolution. Furthermore, a better grasp of the baryon-dark matter link is critical to future cosmological analyses, which are currently impeded by our limited knowledge of baryonic feedback. Here, we measure the kinematic Sunyaev-Zel’dovich (kSZ) effect from the Atacama Cosmology Telescope (ACT), stacked on the luminous red galaxy (LRG) sample of the Dark Energy Spectroscopic Instrument (DESI) imaging survey. This is the first analysis to use photometric redshifts for reconstructing galaxy velocities. Due to the large number of galaxies comprising the DESI imaging survey, this is the highest signal-to-noise stacked kSZ measurement to date: we detect the signal at 13, finding strong evidence that the gas is more spread out than the dark matter, as well as a preference for larger feedback compared to some commonly used state-of-the-art hydrodynamical simulations. Our work opens up the possibility of recalibrating large hydrodynamical simulations using the kSZ effect. In addition, our findings point towards a way of alleviating inconsistencies between weak lensing surveys and cosmic microwave background (CMB) experiments, such as the ‘low ’ tension, and shed light on long-standing enigmas in astrophysics, such as the ‘missing baryon’ problem.
@@ I Introduction
In this work, we measure the baryon profiles around Dark Energy Spectroscopic Instrument (DESI) luminous red galaxies (LRGs) using the stacking of the kSZ effect. The kSZ effect has been measured using a variety of techniques over the years, including the first detections using the pairwise estimator [9, 10], through ‘velocity reconstruction’ templates (and the closely related ‘velocity-inversion’) [11, 12, 13, 14, 15], cross-correlations with redshift fluctuation maps [16], and projected fields techniques [17, 18], among others. Our findings shed light on the complex relationship between baryonic matter and dark matter. The novelty of this work lies in the following three aspects: 1) its pioneering use of photometric redshifts for probing the stacked kSZ effect, which departs from the traditional use of spectroscopic redshifts; 2) challenging the predictions of state-of-the-art hydrodynamical simulations – we find that the feedback processes appear to be stronger in the real Universe than in simulations – in other words, the baryons are expelled more violently from the cores of galaxy groups and clusters; 3) its implications for alleviating or resolving discrepancies in weak lensing analysis; most notably, the ‘low ’ tension found when comparing the inferred values of the clumpiness parameter, , from weak lensing with CMB probes (e.g., see [19] and references therein), and the ‘Lensing is low’ tension, which constitutes a discrepancy between galaxy clustering and galaxy-galaxy lensing [20, 21].
@@ II.1 DESI imaging survey
The Dark Energy Spectroscopic Instrument (DESI) is a robotic, fiber-fed, highly multiplexed spectroscopic surveyor that operates on the Mayall 4-meter telescope at Kitt Peak National Observatory [22]. DESI, which can obtain simultaneous spectra of almost 5000 objects over a field [23, 24, 25], is currently conducting a five-year dark energy survey of about a third of the sky [26]. This campaign will obtain spectra for approximately 40 million galaxies and quasars [27].
In this paper, we use the photometric sample of DESI Luminous Red Galaxies (LRGs) and a higher-density extended LRG sample, referred to as ‘Main LRGs’ and ‘Extended LRGs’ [28, 29, 30, 31, 32]. The parent imaging is the DESI Legacy Imaging Survey, which is a combination of optical and mid-infrared imaging used for DESI target selection from three telescopes: the Dark Energy Camera Legacy Survey (DECaLS), the Mayall -band Legacy Survey (MzLS), and the Beijing–Arizona Sky Survey (BASS). We also make use of the reliable photometric redshifts presented in Zhou et al. [32] for Data Release 9 (DR9) and Data Release 10 (DR10), which have been calibrated via DESI spectroscopic redshifts and thus have only a small fraction of contaminants and a mean error of . For DR10, we adopt the -band estimates (Z_PHOT_MEDIAN_I), as they are more accurate. Similarly to Zhou et al. [32], we define four tomographic bins: [0.4, 0.54], [0.54 0.713], [0.713, 0.86] and [0.86 1.024] (see Table 1 for mean bin redshifts and number of galaxies in the overlap with ACT). We take into account the estimated error in the photometric redshifts for each galaxy, by imposing a maximum cut on the estimated error of . This removes 3% of the galaxies, including outliers with anomalously large redshift errors that could otherwise bias the reconstructed velocities by diluting the density signal [see 33, for tests on simulations]. In App. G, we investigate the effects of not imposing such a cut and find that, for the most part, they are negligible.
@@ II.2 ACT temperature maps
This paper employs the CMB maps maps from the Data Release 6 (DR6) of the Atacama Cosmology Telescope (ACT) [34], a 6 m telescope that was located in the Atacama Desert of Chile and measured the CMB from 2007 to 2022. The multifrequency observational program of DR6 targeted the ‘wide’ field; for this work, we use only the night-time portion of the data taken in the first five observing seasons (2017-2021). In particular, we make use of the harmonic-space Internal Linear Combination (hILC) maps [35], which are constructed by combining the high-resolution (1.5 arcmin) DR6 maps multifrequency observations from 2017 to 2022 at three frequency bands: f090 (77–112 GHz), f150 (124–172 GHz), and f220 (182–277 GHz), alongside Planck data ( 5 arcmin resolution) on large scales [36, 37]. The resulting hILC map covers roughly a third of the sky (32%), with a resolution (i.e., effective beam size Full-Width at Half Maximum) of 1.6 arcmin and mean white noise level in the temperature of approximately 15 K-arcmin. The ACT maps are produced in the plate-carrée (CAR) projection scheme. This analysis uses the first version of the ACT DR6 maps, dr6.01. Similarly to [12], we apply a mask on the ACT map that removes all detected clusters [38] and point sources as well as outliers in any of the filtered temperature bins, since those can bias the stacked profiles.
@@ III.1 Reconstruction
In order to obtain an estimate of the peculiar velocity of each galaxy, we adopt the standard BAO reconstruction method of [40], implemented in the package pyrecon 111https://github.com/cosmodesi/pyrecon [41]. This method yields an estimate of the first-order galaxy displacement field, which can be converted into an estimate for the velocity. A study performed on realistic DESI-like light cone mocks [33] informs us that the cross-correlation coefficient, , between the reconstructed galaxy velocities along the line-of-sight (denoted by subscript ) and the host halo velocities, which captures imperfections in the velocity reconstruction, is about (with an uncertainty of about 10%), the value we adopt in this study, while that of the spectroscopic sample is about . In other words, the reconstruction performance deteriorates by about half when using photometric redshifts, a prediction also confirmed by a similar study in the ‘snapshot geometry’ [42]. Nonetheless, the measurement around the photometric LRG sample yields a high signal-to-noise ratio (SNR) due to the large number of galaxies comprising the imaging survey. We note that LRGs are found mostly in galaxy groups with a mean linear bias of [43]. In addition to the cuts in [12], for our fiducial analysis, we remove outliers in the reconstructed velocities: in particular, we ensure that the number of galaxies in each velocity bin is symmetric around the mean by random downsampling, which avoids unwanted biases from massive clusters and guarantees that in the absence of a kSZ signal, our estimator yields zero mean signal. In App. G, we show that these corrections have a mostly negligible effect: the uncorrected measurements tend to have marginally higher SNR (as a result of the larger number of galaxies in the samples) except for bin 3, which has a lower SNR due to its slightly asymmetric line-of-sight velocity distribution that improves once we remove the outliers.
@@ III.2 Estimator
Once we have estimates of the velocities, we can measure the CMB temperatures around each galaxy using a compensated aperture photometry (CAP) filter [12]. There are several benefits to using the CAP filter. Unlike a matched filter, the CAP filter is agnostic about the profile shape and because it is measured at different radii, it allows us to reconstruct the spherically averaged profiles. It is very effective at removing the large-scale primary CMB, as well as the uncorrelated part of the signal, leaving, in principle, only the sum of the one- and two-halo terms [see 44, for tests on simulations]. Additionally, the CAP-filtered profiles behave similarly to a cumulative density profile for large radii. The CAP filter is defined as:
@@
where is the rms of the radial component of the reconstructed velocities, , is the speed of light, and is the cross-correlation coefficient between the reconstructed and true velocity, evaluated from mock simulations. We use the publicly available pipeline ThumbStack222https://github.com/EmmanuelSchaan/ThumbStack to apply the estimator to the DESI and ACT data.
To interpret the kSZ profiles obtained through the above method, we need to estimate the covariance of the stacked measurements. We do so by applying a bootstrap resampling to the signal at each galaxy location. In particular, we draw 10,000 realizations of the galaxy catalogs (with repetition), and infer the covariance matrices from the scatter across the resampled stacked profiles (see Section B).
@@ IV.1 Stacked kSZ signal per redshift bin
| LRG sample | # of galaxies | |||
| Extended DR9 | 963,631 | 0.47 | 43.2 | 6.4 |
| Extended DR9 | 1,658,313 | 0.63 | 69.4 | 7.1 |
| Extended DR10 | 1,951,646 | 0.79 | 80.3 | 8.2 |
| Extended DR10 | 1,690,171 | 0.92 | 34.0 | 4.6 |
| Extended DR9+10 all | 6,850,072 | 0.75 | 203.0 | 13.5 |
| Main DR9 | 422,350 | 0.47 | 25.2 | 4.4 |
| Main DR9 | 795,393 | 0.63 | 78.8 | 7.8 |
| Main DR10 | 753,945 | 0.79 | 65.9 | 7.3 |
| Main DR10 | 629,367 | 0.93 | 20.8 | 3.2 |
| Main DR9+10 all | 2,882,904 | 0.74 | 166.7 | 12.1 |
Fig. 1 shows the stacked kSZ signal (Eq. 4) from ACT for the four photometric bins of the DR10 Extended and Main LRG samples obtained from the DESI Imaging Survey (see Section II.1). The signal is shown as a function of CAP filter radius and detected at a high significance. We note that the error bars at large aperture are highly correlated (see App. B on the covariance estimation). To aid interpretation, we convert the and axis into proper distance and CAP-filtered optical depth (), respectively, calculated at the mean redshift, . We note that the optical depth measures the integrated gas density along the line-of-sight,
@@
In Table 1, we show the significance of detection and corresponding SNR and chi-squared values for each of the four bins as well as the combination of all four bins. We define the SNR with respect to null333Corresponding to the inverse of the fractional error on the fit of a single amplitude parameter to the data. This often called the “detection significance” in number of Gaussian standard deviations . as , with best-fit coming from the Illustris-1 simulation444https://www.tng-project.org/data/ curve at (see Fig. 2 for more details) with a free parameter for the amplitude. We detect the signal at 13 as shown in Table 1.
Here we have defined the metric in the standard way using the covariance matrix (see App. B):
where is our stacked kSZ measurement (data vector), and is the model we are comparing against (if null, ).
To take advantage of the larger number of objects in DR9 and their relatively small photometric errors at low (compared with high ), we quote the DR9 results for bin 1 and 2 and the DR10 results for bin 3 and 4. For the combined ‘DR9+10 all,’ we use DR10 reconstructed velocities where available and otherwise DR9 ones in order to maximize the signal-to-noise. We provide detailed comparison between DR9 and DR10 in Tables 2 and 3 finding a high level of congruence between the two. Compared with the previous measurement using CMASS LRGs and ACT DR4 [12], (same number of radial bins), we see that the total chi-squared for our Extended and Main samples is and 170, respectively, or about twice higher. In App. F, we find our curves to be in excellent agreement with [12].
Intuitively, one can think of these curves as showing roughly the cumulative gas density distribution. At large radii, we expect the profiles to become shallower as more and more of the gas is enclosed. The gas profiles are steep beyond the virial radius, suggesting that a significant fraction of the gas resides beyond it (see discussion of Fig. 2, where we test this conjecture). The decline in amplitude at large apertures of bin 3 is due to noise from the primary CMB and the larger fraction of photometric outliers in that bin (see App. G).
@@ IV.2 Comparison with simulations
We first compare our data to dark matter simulations: the dark matter profiles are obtained by stacking on dark-matter-only maps computed using the TNG300-1-Dark simulation (dark-matter-only counterpart to TNG300-1), with the results shown in Fig. 2.
Quantifying the level of agreement or disagreement between data and simulations without the use of external measurements or priors is challenging because the amplitude of the results is very sensitive to the mass of the sample and to finite-box effects in simulations. The same selection applied to different simulations will result in different mean masses, complicating the comparison.
Similarly to [45], we adopt two methods, each with advantages and drawbacks. In the first, we normalize all simulations to contain the same gas mass at large aperture (5 arcmin in this case), matching the value of the best-fitting (lowest with the data) simulation, in this case Illustris-1. This ensures that we are comparing halos of the same mass across all simulations, and that such gas mass is compatible with the data. We call this “fixed amplitude”. Then we quantify the relative difference with respect to the best-fit simulation as , where the DM amplitude has been fixed as described. We find and 43, for the Main and Extended samples, respectively. We note that a major drawback of this procedure is that, while it ensures consistency of gas mass between the data and the simulations, it doesn’t properly marginalize over the uncertainty in this quantity, resulting in larger values of .
Alternatively, we can let the amplitude of each simulation be completely free and fit it to the data: this fully marginalizes over the uncertainty in all of the data points. The drawback is that it is allowed to find best-fit solutions with very low gas or halo mass, potentially ruled out by external data (for example, weak lensing or fits to the galaxy correlation function). Nonetheless, this is the most conservative comparison we can perform without joint analysis of external data, and therefore we’ll take it as the baseline for our conclusions. We can define a “free amplitude” , where the DM curve has the best-fit amplitude. For both samples, we find . Given that even in the more conservative case of “free amplitude”, the dark matter curve is a much poorer fit to the data, we conclude that we find strong evidence that the gas as measured by the kSZ effect, is significantly more extended than the dark matter.
We note that the limitations of both approaches can be mitigated by the inclusion of external data such as weak lensing or galaxy clustering (among others), which independently constrain the mass of the sample [46]. Such joint analyses will be able to sharpen the conclusions significantly, but they are beyond the scope of this paper and will be the subject of future work.
@@
We then study the comparison between the state-of-the-art hydrodynamical simulations IllustrisTNG-300-1 (TNG300-1) [47] and the measured gas profiles from the data. We also show a comparison with the older model of the Illustris-1 simulation [48], which has known shortcomings with predicting observations such as galaxy morphologies and colors, various cluster properties, and gas fractions [49, 50, 51, 47]. We provide details on how the measurements on simulations are made in App. C. In App. E, we split the galaxies into stellar mass bins estimated in [31] and detect the mass evolution of the signal at high significance.
@@
We perform a number of tests on both hydrodynamical simulations to ensure that the gap we see is significant: we vary the satellite fraction from 0% to 30%555Centrals are selected as the most massive subhalo in each host halo. To vary the satellite fraction, we randomly discard centrals or satellites. (note that the constraints from [43] are ), we put all galaxies at the centers of their host halos, we vary the number density (between half and twice the fiducial value, i.e. from to ) and thus stellar mass threshold of the extracted LRGs, and we add noise to the halo velocities. In Fig. 2, we show the default scenario for both simulations, and in shaded color for TNG300-1, we display the minimum and maximum deviations from the default caused by all the aforementioned alternative scenarios. Despite considering scenarios that push the bounds of physically reasonable models, we see that the gap between simulations and observations persists.
To quantify this, we can once again define a “fixed amplitude” and “free amplitude” , with the best-fitting curve as usual being Illustris-1. We find and 19 for the Main and Extended samples, and for both samples. Even in the very conservative “free amplitude” case, we find a preference for a more extended profile and hence larger feedback compared to TNG300-1.
IV.3 Implications for lensing
Given the fact that there appears to be a mismatch between the baryons and the dark matter distributions within a few virial radii, and that baryons amount for 16% of the mass, our findings imply that baryons have a non-negligible effect on lensing, especially on small scales. While we leave a detailed study to future work, we note that with a similar baryon distribution around the BOSS CMASS galaxies as measured by kSZ [12], the authors were able to show a 15% suppression on the lensing signal on scales smaller than 1 Mpc, and a few percent suppression up to several Mpc [52]. This is consistent with evacuating the majority of the baryons from sub-Mpc scales, leading to a suppression in the lensing signal of order the baryon mass fraction, and significantly mitigating the “Lensing is low” [20] problem. We expect similar behavior here.
On larger scales, one may wonder if baryons play a part in the “ tension” [19] as measured by several weak lensing surveys: the impact will likely depend on the analysis choices of each specific experiment and their scale cuts, which are typically determined using lower-feedback simulations. For example, a joint analyses of DES data with the BOSS kSZ measurements hint that baryon effects can mitigate part of the tension [53, 46]. Given that the measurements presented here are twice as precise as those used in previous work and cover a broader redshift range, similar analyses should be able to significantly strengthen the conclusions. We leave a detailed study to upcoming work.